# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 4 Transfer reactions with weakly bound nuclei
To see how this method works in practice, let us consider as a particular case a process of the form $A(d,p)B$, schematically depicted in Fig. 21. Using the post-form representation, the transition matrix for this process can be written as
@@
where $I$ is the spin of $A$, $l$ and $j$ the orbital and total ($\vec{j}=\vec{\ell}+s$) angular momentum of the valence particle and $\phi_{\ell j}({\vec{r}}_{nA})$ is a function describing the neutron-core relative motion. The normalization
@@
can be regarded as a spectroscopic factor for the configuration $\{\ell,j\}$. [^2] Strictly, the functions $\phi_{lj}$ and the spectroscopic factors ${\cal S}_{\ell j}$ depend in general on the core state $I$ but, in the present case, this quantum number can be readily inferred from the $l,j$ values, so it is omitted for brevity.
If the interactions $U_{pA}$ and $U_{pB}$ are assumed to be independent of the target degrees of freedom ($\xi_{A}$), the integral in $\xi_{A}$ can be readily performed, transforming (40) into
@@
In DWBA, the three-body wavefunction $\Psi^{3b,(+)}_{d}$ is further approximated by $\Psi^{\mathrm{(+)}}\approx\chi^{(+)}_{d}({\vec{R}})\varphi_{d}({\vec{r}})$. When the transfer reaction involves weakly-bound nuclei, including the deuteron discussed here, this choice is not well justified. Due to the presence of the short-range $V_{pn}$ interaction, the evaluation of DWBA matrix element is mostly sensitive to small $p$-$n$ separations. These configurations do not only contain contributions coming from the deuteron ground state, but also from $p$-$n$ unbound states. By contrast, the deuteron optical potential describing deuteron-target elastic scattering is not restricted to small $p$-$n$ distances. Therefore, the use of the deuteron optical potential in the $(d,p)$ or $(p,d)$ transfer amplitude is likely to lack important deuteron breakup components. This problem was recognized long ago and several solutions have been proposed. One of the first and most popular ones is the adiabatic method of Johnson and Soper Ron70 .
@@
We see that, in this limit, the adiabatic theory of the transfer amplitude adopts a form akin to that found in DWBA, but this analogy is only formal because the function $\chi_{JS}^{(+)}({\vec{R}})$ includes contributions from breakup and the potential $U_{JS}(R)$ may have little to do with the optical potential describing the deuteron elastic scattering. Due to the adiabatic approximation, the JS theory is not expected to be accurate at low incident energies.
The adiabatic approximation is equivalent to neglecting the excitation energy of the states of the projectile Ron70 . The adiabatic wave function takes into account the excitation to breakup channels, assuming that these states are degenerate in energy with the projectile ground state, as illustrated in Fig. 22(b). Therefore, the ADWA approach takes into account, approximately, the effect of deuteron break-up on the transfer cross section, within the adiabatic approximation. So, it should be well suited to describe deuteron scattering at high energies, around 100 MeV per nucleon. Systematic studies Har71 ; Sat71 ; Wal76 have shown that ADWA is superior to standard DWBA for $(d,p)$ scattering at these relatively high energies.
@@
Although the zero-range adiabatic model of JS provides a systematic improvement over the conventional DWBA, there are situations in which the former fails to reproduce the experimental data Ste86 ; Ron89 . Models which go beyond the zero-range and adiabatic approximations are therefore needed. One of such models is the Weinberg expansion method of Johnson and Tandy (JT) JT74 . The idea is to expand $\Psi^{\mathrm{(+)}}$ in terms of a set of functions which are complete within the range of $V_{pn}$. A convenient choice is the set of Weinberg states (also called Sturmians), given by
$\Psi^{(+)}(\vec{R},\vec{r})=\sum_{i=0}^{N}\phi^{W}_{i}(\vec{r})\chi^{W}_{i}(\vec{R})\,,$ (47)
@@
where $\varepsilon_{d}=2.225$ MeV is the deuteron binding energy and where $\alpha_{i}$ are the eigenvalues, to be determined along with the eigenfunctions. Beyond the range of the potential all the Weinberg states decay exponentially, like the deuteron ground-state wave function. For $i=0$, $\alpha_{0}=1$ and so $\phi^{W}_{0}({\vec{r}})$ is just proportional to the deuteron ground state. As $i$ and $\alpha_{i}$ increase, they oscillate more and more rapidly at short distances. The Weinberg states form a complete set of functions of ${\vec{r}}$ within the range of the potential $V_{pn}$. They are well suited to expand $\Psi^{(+)}$ in this region, as it is required by the amplitude (40). They do not satisfy the usual orthonormality relation but the less conventional one
@@
If we retain in (47) only the leading term, $\Psi^{(+)}(\vec{R},\vec{r})\approx\phi^{W}_{0}\chi^{W}_{0}(\vec{R})$, one finds JT74 that $\chi^{W}_{0}$ verifies the single-channel equation
@@
$U_{JT}(R)=\frac{\langle\phi_{pn}(\vec{r})|V_{pn}(U_{nA}+U_{pA})|\phi_{pn}(\vec{r})\rangle}{\langle\phi_{pn}(\vec{r})|V_{pn}|\phi_{pn}(\vec{r})\rangle}.$ (51)
@@
In Fig. 23 we present a comparison between the DWBA and ADWA methods for the 56Ni(p,d)57Ni (left) and 48Ca(d,p)49Ca reactions. For the ADWA calculations, the prescriptions of JS and JT are shown. It is clearly seen that the ADWA, in both its zero-range and finite-range forms, provides an improved description of the data as compared to the DWBA method.
Figure 23: Comparison of the ADWA and DWBA methods for the 56Ni($p$,$d$)57Ni (left) and 48Ca($d$,$p$)49Ca reactions at incident energies of 37 MeV and 10 MeV, respectively. The nucleon optical potentials are taken from the global CH89 systematics while Daenick global potential is used for the deuteron optical potential in the DWBA calculations. For the ADWA calculations, the prescriptions of Johnson-Soper (JS) and Johnson-Tandy (JT) are shown. Quoted from Ref. Tim20 .
An appealing feature of the ADWA method is that its ingredients are completely determined by experiments. These ingredients are the proton-target and neutron-target optical potentials, evaluated at half of the deuteron incident energy, as well as the well-known proton-neutron interaction. On the negative side, the ADWA approach does not consistently describe elastic scattering and nucleon transfer. Although, physically, elastic scattering, transfer and break-up should be closely related by flux conservation, this connection is not present in ADWA. Furthermore, the arguments leading to ADWA are strongly dependent on the assumption that the transfer is governed by a short-range operator. Thus, it is not obvious that the approximations remain valid for other weakly bound systems, like 11Be. Even in the case of $(d,p)$ scattering, the transfer matrix element is determined, in addition to the $n$-$p$ interaction, by the proton-target and proton-composite interactions which define the remnant term. The role of these terms, which would have contributions of three-body configurations in which proton and neutron are not so close together, is not clear a priori. The latter problem can be avoided by using an alternative expression for the scattering amplitude. Following Goldberg and Watson Gol64 , one can choose the auxiliary interaction $U_{pB}$ that appears in the remnant term of Eq. (44) to cancel exactly the core-target interaction, $U_{pA}$. This results in an alternative, but still exact, scattering amplitude (the SF is omitted for simplicity), namely,
@@
In Tim99 Timofeyuk and Johnson use an adiabatic approximation for $\tilde{\Phi}^{(-)}_{pB}$ to produce a tractable expression that still includes recoil excitation and breakup effects (REB), while keeping the matrix element constrained to the range of the $V_{pn}$ interaction:
@@
In this expression, $\boldsymbol{k}_{\alpha}$ is the relative momentum between $B$ and $p$ and $\alpha=m_{n}/m_{B}$, and it was found to produce a good agreement between this calculation and experimental data for the 16O$(d,p)^{17}$ and 10Be$(d,p)^{11}$e reactions. Nonetheless, given that this expression is better suited for the transfer of weakly-bound and halo nuclei, the ADWA model has enjoyed more widespread use.
@@
Another way of accounting for the breakup channels in transfer reactions is by insertion of the CDCC wavefunction, Eq. (8), in the transition amplitude (44). The resultant amplitude is formally analogous to that found in the CCBA method Sat83 , so we refer to it as CDCC-BA approximation. This is expected to be a good approximation since the CDCC expansion is accurate for small $p$-$n$ separations, for which the matrix transition amplitude presents the largest magnitude. The resultant amplitude, however, will be clearly much more computationally demanding than that obtained with the DWBA or ADWA methods.
In Fig. 24 we compare the performance of the DWBA, ADWA and CDCC-BA methods for the 58Ni($d$,$p$)59Ni reaction at $E_{d}=10$ MeV (upper panel) and $E_{d}=56$ MeV (lower panel), taken from Ref. Pan14 . It becomes clear that the ADWA and CDCC methods provide a more reliable description of the data compared to the traditional DWBA approach. In Cha17 , a systematic comparison between the CDCC-BA method and an adiabatic approximation (CDCC-AD) was performed for $(d,p)$ reactions on various targets, showing good agreement between both, except in some cases, such as when the deuteron beam energy is small and similar to the binding energy of the nucleon in the target or when the binding energy of the nucleon is very small ($S_{n}~{}\simeq 0.1$ MeV). It should be noted that in this work the adiabatic approximation is performed by setting the nominal energies of the continuum bins to that of the bound state, which is formally equivalent to the ADWA approach, although the calculation is rather different, so extensions of these results to standard ADWA calculations should be made with caution.
@@ 4.1 Transfer reactions populating unbound systems
So far, we have considered transfer reactions as a tool for investigating the bound states of a given nucleus. However, in a rearrangement process, the transferred particle can also populate unbound states of the final nucleus. This opens the possibility of studying and characterizing structures in the continuum, such as resonances or virtual states. In fact, since weakly-bound nuclei usually have almost no excited bound states and their binding energy is low, the matching conditions for transfer reactions Bri72 favours the population of highly excited states of the residual nucleus, including continuum states.
@@
An example is shown in Fig. 26, which corresponds to the differential cross section, as a function of the $n$-9Li relative energy, for the reaction 2H(9Li,$p$)10Li at 2.36 MeV/u (upper) and 11.1 MeV/u (lower panel), corresponding to the measurements of Refs. Jep06 and TRIUMF Cav17 , respectively. The lines are the results of transfer-to-the-continuum calculations populating 10Li continuum states using the same 10Li structure in both cases. The figure shows the separate contribution of the s-wave ($1^{-}$, $2^{-}$) and p-wave ($1^{+}$, $2^{+}$) continuum states. The strength of the measured cross section close to zero energy is due to the presence of a virtual state in the 10Li continuum, whereas the peak around 0.4 MeV is due to a $p_{1/2}$ resonance. This is an example of how the use of transfer reactions can provide information on the continuum structure of weakly-bound or even unbound systems. For more details on the calculations, see Ref. Mor19 .
@@ 4.2 Transfer reactions involving three-body projectiles
The formalism presented in the previous sections can be extended and applied to more complex systems. We discuss here the interesting case of the stripping reaction induced by a three-body Borromean system on a proton target, in which one of the projectile fragments is transferred to the target, leaving a two-body residual unbound system. To be more specific, we consider the case of a two-neutron halo projectile, such as 11Li, 6He. In these cases, given the unstable nature of the projectile, the reaction experiment must be performed in inverse kinematics. This process can be schematically represented as (see Fig. 27).
@@
$\mathcal{T}_{if}=2~{}\langle\Psi_{f}^{(-)}(\vec{x},\vec{R}^{\prime},\vec{r}^{\prime})|V_{pN_{1}}+U_{pB}-U_{pA}|\Phi_{A}(\vec{x},\vec{y})\chi_{pA}^{(+)}(\vec{R})\rangle,$ (58)
where $\Phi_{A}$ represents the ground-state wave function of the initial three-body composite, $\chi^{(+)}_{pA}$ is the distorted wave generated by the auxiliary potential $U_{pA}$, and $\Psi^{(-)}_{f}$ is the exact four-body wave function for the outgoing $d$-$B$ system. Notice the explicit factor 2 coming from the two identical neutrons in the initial wave function.
@@
In writing the transition amplitude in the form (58), one implicitly uses a participant/spectator approximation, assuming that the reaction occurs due to the interaction of the incident proton with a single neutron $(N_{1})$ of $A$ (the participant), leaving the residual $B=N_{2}+C$ system unperturbed.
To reduce (58) to a tractable form, one can approximate the exact wave function $\Psi^{(-)}_{f}$ by the factorized form,
@@
After integration over the angle of $\vec{q}$ and transforming from momentum to energy, expression (62) provides double differential cross section with respect to the scattering angle and the excitation energy of the residual $N_{2}+C$ system. This observable can in principle be reconstructed experimentally from the measured energy and angle of the outgoing deuteron. To illustrate this formalism, we show in Fig. 28 results for the reaction 11Li(p,d)10Li Cas17 , recently measured at TRIUMF by Sanetullaev et al. San16 .
@@
For the calculations presented, it was assumed that the two halo neutrons of 11Li have the same angular momentum $\vec{J}=\vec{L}+\vec{s}_{2}$. This condition is strictly fulfilled when the spin of the core is neglected and is also a good approximation when using $I=3/2^{-}$. Only the configurations $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ were considered, as other contributions, while present in the three-body calculations, give negligible components in the ground state of 11Li. The considered 9Li+n potential contains terms that couple the spin of the remaining neutron and 9Li in 10Li leading to splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ states and of the virtual state in the $s$-wave in $1^{-}$ and $2^{-}$ components. These components are shown in the top panel of Fig. 28, which depicts the relative energy spectrum between $n$ and 9Li. In red, the $p-$wave components are presented, where the splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ resonances at 0.37 and 0.61 MeV is evident, while in green the $s$-wave conributions are shown, presenting a virtual state (with scattering length $a=-37.9$ fm) only in the $2^{-}$ component. The total is shown in dashed blue, while its folding with the experimental resolution given in Ref. San16 is presented in the solid blue line. It is rather remarkable how the low-energy peaks are smoothed out by the experimental resolution leading to a relatively featureless distribution, which could explain why the fine details of 10Li are not observed in this experiment. In the bottom panel, the energy-integrated angular distribution of the outgoing deuteron is presented. It has been assumed that the subtraction of the non-resonant background performed in the analysis of the data completely excludes the $s$-wave component, so experimental data are compared only with the cross section corresponding to neutrons transferred from the $p_{1/2}$ orbitals, finding an excellent agreement both in shape and magnitude. Note that the experimental data have been measured at angles for which the $s_{1/2}$ contribution is minimal. This may explain why in our energy distribution the contribution from $s$ waves is quite important at low energies, while experimentally there is no direct sign of this effect.
Using the methods described in Sec. 4.1 to consider continuum states of the $p-N$ subsystem, a similar formalism can be used to study $(p,pN)$ reactions on three-body nuclei, simply considering the auxiliary amplitude (61) for continuum $p-N$ states, which are typically chosen to have a definite $p-N$ energy, angular momentum and parity:
@@
and where the outgoing $\chi_{pNB}^{(-)}$ functions can be computed using a CDCC solution for the final 3-body system Mor15 . Figures 29 shows results for the $n$-${}^{9}\mathrm{Li}$ relative energy distribution in the 11Li$(p,pn)^{10}$i reaction at 280 MeV/nucleon, using the same model for 11Li and 10Li as for the $(p,d)$ reaction in Fig. 28. The contributions of the different components are shown on the top panel, where again a splitting of the $p-$wave resonance is visible, as well as the virtual state only in the $2^{-}$ component, the main difference to the $(d,p)$ case being the increased importance of the $s$-wave components. In the bottom panel, the theoretical distribution is presented convoluted with the experimental resolution and compared with the experimental data, showing an excellent agreement with it, although the experimental resolution again significantly smooths the distribution.
It is a success of the model model that it is able to reproduce experimental data at $\sim$6 MeV per nucleon (the $(p,d)$ reaction) and at $\sim$280 MeV per nucleon (the $(p,pN)$ reaction), although perhaps not unexpected, given the similarities of both reactions when viewed as transfer to bound and continuum states of the $p-n$ system.

### v2

@@ 4 Transfer reactions with weakly bound nuclei
To see how this method works in practice, let us consider as a particular case a process of the form $A(d,p)B$, schematically depicted in Fig. 25. Using the post-form representation, the transition matrix for this process can be written as Sat83
@@
where $I$ is the spin of $A$, $\ell$ and $j$ the orbital and total ($\vec{j}=\vec{\ell}+s$) angular momentum of the valence particle and $\phi_{\ell j}({\vec{r}}_{nA})$ is a function describing the neutron-core relative motion. The normalization
@@
can be regarded as a spectroscopic factor for the configuration $\{\ell,j\}$. [^2] Strictly, the functions $\phi_{\ell j}$ and the spectroscopic factors ${\cal S}_{\ell j}$ depend in general on the core state $I$ but, in the present case, this quantum number can be readily inferred from the $\ell,j$ values, so it is omitted for brevity.
If the interactions $U_{pA}$ and $U_{pB}$ are assumed to be independent of the target degrees of freedom ($\xi_{A}$), the integral in $\xi_{A}$ can be readily performed, transforming Eq. (48) into
@@
In DWBA, the three-body wavefunction $\Psi^{3b,(+)}_{d}$ is further approximated by $\Psi^{\mathrm{(+)}}_{d}\approx\chi^{(+)}_{d}({\vec{R}})\varphi_{d}({\vec{r}})$. When the transfer reaction involves weakly-bound nuclei, including the deuteron discussed here, this choice is not well justified. Due to the presence of the short-range $V_{pn}$ interaction, the evaluation of DWBA matrix element is mostly sensitive to small $p$-$n$ separations. These configurations do not only contain contributions coming from the deuteron ground state, but also from $p$-$n$ unbound states. By contrast, the deuteron optical potential describing deuteron-target elastic scattering is not restricted to small $p$-$n$ distances. Therefore, the use of the deuteron optical potential in the $(d,p)$ or $(p,d)$ transfer amplitude is likely to lack important deuteron breakup components. This problem was recognized long ago and several solutions have been proposed. One of the first and most popular ones is the adiabatic method of Johnson and Soper Ron70 .
@@
We see that, in this limit, the adiabatic theory of the transfer amplitude adopts a form akin to that found in DWBA, but this analogy is only formal because the function $\chi_{JS}^{(+)}({\vec{R}})$ includes contributions from breakup and the potential $U_{JS}(R)$ may have little to do with the optical potential describing the deuteron elastic scattering, so $\chi_{JS}^{(+)}$ does not provide a good description of the elastic channel. Due to the adiabatic approximation, the JS theory is not expected to be accurate at low incident energies.
The adiabatic approximation is equivalent to neglecting the excitation energy of the projectile states, Ron70 , which amounts at setting $H_{bx}\to\varepsilon_{0}$ in the three-body Hamiltonian (6). The adiabatic wave function takes into account the excitation to breakup channels, assuming that these states are degenerate in energy with the projectile ground state, as illustrated in Fig. 26(b). Therefore, the ADWA approach takes into account, approximately, the effect of deuteron break-up on the transfer cross section, within the adiabatic approximation. So, it should be well suited to describe deuteron scattering at high energies, around 100 MeV per nucleon. Systematic studies Har71 ; Sat71 ; Wal76 have shown that ADWA is superior to standard DWBA for $(d,p)$ scattering at these relatively high energies.
@@
Although the zero-range adiabatic model of JS provides a systematic improvement over the conventional DWBA, there are situations in which the former fails to reproduce the experimental data Ste86 ; Ron89 . Models which go beyond the zero-range and adiabatic approximations are therefore needed. One of such models is the Weinberg expansion method of Johnson and Tandy (JT) JT74 . The idea is to expand $\Psi_{d}^{\mathrm{3b,(+)}}$ in terms of a set of functions which are complete within the range of $V_{pn}$. A convenient choice is the set of Weinberg states (also called Sturmians), given by
$\Psi_{d}^{3b,(+)}(\vec{R},\vec{r})=\sum_{i=0}^{N}\phi^{W}_{i}(\vec{r})\chi^{W}_{i}(\vec{R})\,,$ (55)
@@
where $\varepsilon_{d}=2.225$ MeV is the deuteron binding energy and where $\alpha_{i}$ are the eigenvalues, to be determined along with the eigenfunctions. Beyond the range of the potential all the Weinberg states decay exponentially, like the deuteron ground-state wave function. For $i=0$, $\alpha_{0}=1$ and so $\phi^{W}_{0}({\vec{r}})$ is just proportional to the deuteron ground state. As $i$ and $\alpha_{i}$ increase, they oscillate more and more rapidly at short distances. The Weinberg states form a complete set of functions of ${\vec{r}}$ within the range of the potential $V_{pn}$. They are well suited to expand $\Psi_{d}^{3b,(+)}$ in this region, as is required by the amplitude in Eq. (48). They do not satisfy the usual orthonormality relation but the less conventional one
@@
If we retain in Eq. (55) only the leading term, $\Psi_{d}^{3b,(+)}(\vec{R},\vec{r})\approx\phi^{W}_{0}\chi^{W}_{0}(\vec{R})$, one finds JT74 that $\chi^{W}_{0}$ verifies the single-channel equation
@@
$U_{JT}(R)=\frac{\langle\phi^{W}_{0}(\vec{r})|V_{pn}(U_{nA}+U_{pA})|\phi^{W}_{0}(\vec{r})\rangle}{\langle\phi^{W}_{0}(\vec{r})|V_{pn}|\phi^{W}_{0}(\vec{r})\rangle}.$ (59)
@@
In Fig. 27 we present a comparison between the DWBA and ADWA methods for the 56Ni(p,d)57Ni (left) and 48Ca(d,p)49Ca (right) reactions. For the ADWA calculations, the prescriptions of JS and JT are shown. It is clearly seen that the ADWA, in both its zero-range and finite-range forms, provides an improved description of the data as compared to the DWBA method.
Figure 27: Comparison of the ADWA and DWBA methods for the 56Ni($p$,$d$)57Ni (left) and 48Ca($d$,$p$)49Ca (right) reactions at incident energies of 37 MeV and 10 MeV, respectively. The nucleon optical potentials are taken from the global CH89 CH89 systematics while the Daehnick DCV80 global potential is used for the deuteron optical potential in the DWBA calculations. For the ADWA calculations, the prescriptions of Johnson-Soper (JS) and Johnson-Tandy (JT) are shown. Quoted from Ref. Tim20 .
An appealing feature of the ADWA method is that its ingredients are completely determined by experiments. These ingredients are the proton-target and neutron-target optical potentials, evaluated at half the deuteron incident energy, as well as the well-known proton-neutron interaction. On the negative side, the ADWA approach does not consistently describe elastic scattering and nucleon transfer. Although, physically, elastic scattering, transfer and break-up should be closely related by flux conservation, this connection is not present in ADWA. Furthermore, the arguments leading to ADWA are strongly dependent on the assumption that the transfer process is governed by a short-range operator. Thus, it is not obvious that the approximations remain valid for other weakly bound systems, like 11Be. Even in the case of $(d,p)$ scattering, the transfer matrix element is determined, in addition to the $n$-$p$ interaction, by the proton-target and proton-composite interactions which define the remnant term. The role of these terms, which would have contributions of three-body configurations in which proton and neutron are not so close together, is not clear a priori. The latter problem can be avoided by using an alternative expression for the scattering amplitude. Following Goldberg and Watson Gol64 , one can choose the auxiliary interaction $U_{pB}$ that appears in the remnant term of Eq. (52) to cancel exactly the core-target interaction, $U_{pA}$. This results in an alternative, but still exact, scattering amplitude (the spectroscopic factor is omitted for simplicity), namely,
@@
In Tim99 Timofeyuk and Johnson use an adiabatic approximation for $\tilde{\Phi}^{(-)}_{pB}$ to produce a tractable expression that still includes recoil excitation and breakup effects, while keeping the matrix element constrained to the range of the $V_{pn}$ interaction:
@@
In this expression, $\boldsymbol{k}_{\alpha}$ is the relative momentum between $B$ and $p$ and $\alpha=m_{n}/m_{B}$, and it was found to produce a good agreement between this calculation and experimental data for the 16O$(d,p)^{17}$ and 10Be$(d,p)^{11}$e reactions. Nevertheless, given that this expression is better suited for the transfer of weakly bound and halo nuclei, the ADWA model has enjoyed more widespread use.
@@
Another way of accounting for the breakup channels in transfer reactions is by insertion of the CDCC wavefunction, Eq. (9), in the transition amplitude of Eq. (52). The resultant amplitude is formally analogous to that found in the CCBA method Sat83 , so we refer to it as CDCC-BA approximation. This is expected to be a good approximation since the CDCC expansion is accurate for small $p$-$n$ separations, for which the matrix transition amplitude presents the largest magnitude. The resultant amplitude, however, will be clearly much more computationally demanding than that obtained with the DWBA or ADWA methods.
In Fig. 28 we compare the performance of the DWBA, ADWA and CDCC-BA methods for the 58Ni($d$,$p$)59Ni reaction at $E_{d}=10$ MeV (upper panel) and $E_{d}=56$ MeV (lower panel), taken from Ref. Pan14 . It becomes clear that the ADWA and CDCC methods provide a more reliable description of the data compared to the traditional DWBA approach. In Cha17 , a systematic comparison between the CDCC-BA method and an adiabatic approximation (CDCC-AD) was performed for $(d,p)$ reactions on various targets, showing good agreement between both, except in some cases, such as when the deuteron beam energy is small and similar to the binding energy of the nucleon in the target or when the binding energy of the nucleon is very small ($S_{n}~{}\simeq 0.1$ MeV). It should be noted that in that work the adiabatic approximation is performed by setting the nominal energies of the continuum bins to that of the bound state, which is formally equivalent to the ADWA approach, although the calculation is rather different, so extensions of these results to standard ADWA calculations should be made with caution.
@@ 4.1 Transfer reactions populating unbound systems
So far, we have considered transfer reactions as a tool for investigating the bound states of a given nucleus. However, in a rearrangement process, the transferred particle can also populate unbound states of the final nucleus. This opens the possibility of studying and characterizing structures in the continuum, such as resonances or virtual states. In fact, due to the matching conditions for this type of processes Bri72 , reactions induced by weakly-bound nuclei favour the population of highly excited states of the residual nucleus, including those above the breakup threshold.
@@
An example is shown in Fig. 30, which corresponds to the differential cross section, as a function of the $n$-9Li relative energy, for the reaction 2H(9Li,$p$)10Li at 2.36 MeV/u (upper) and 11.1 MeV/u (lower panel), corresponding to the measurements of ISOLDE Jep06 and TRIUMF Cav17 , respectively. The lines are the results of transfer-to-the-continuum calculations populating 10Li continuum states using the same 10Li structure in both cases. The figure shows the separate contribution of the s-wave ($1^{-}$, $2^{-}$) and p-wave ($1^{+}$, $2^{+}$) continuum states. The strength of the measured cross section close to zero energy is due to the presence of a virtual state in the n+9Li $s$-wave, whereas the peak around 0.4 MeV is due to a $p_{1/2}$ resonance. This is an example of how the use of transfer reactions can provide information on the continuum structure of weakly-bound or even unbound systems. For more details on the calculations, see Ref. Mor19 .
@@ 4.2 Transfer reactions involving three-body projectiles
The formalism presented in the previous sections can be extended and applied to more complex systems. We discuss here the interesting case of the stripping reaction induced by a three-body Borromean system, in which one of the projectile fragments is transferred to the target, leaving a two-body residual unbound system. To be more specific, we consider the case of a two-neutron halo projectile, such as 11Li, 6He impinging on a proton target. In these cases, given the unstable nature of the projectile, the reaction experiment must be performed in inverse kinematics. This process can be schematically represented as (see Fig. 31).
@@
$\mathcal{T}_{if}=\sqrt{2}~{}\langle\Psi_{f}^{(-)}(\vec{x},\vec{R}^{\prime},\vec{r}^{\prime})|V_{pN_{1}}+U_{pB}-U_{pA}|\Phi_{A}(\vec{x},\vec{y})\chi_{pA}^{(+)}(\vec{R})\rangle,$ (66)
where $\Phi_{A}$ represents the ground-state wave function of the initial three-body composite, $\chi^{(+)}_{pA}$ is the distorted wave generated by the auxiliary potential $U_{pA}$, and $\Psi^{(-)}_{f}$ is the exact four-body wave function for the outgoing $d$-$B$ system. Notice the explicit factor $\sqrt{2}$ coming from the two identical neutrons in the initial wave function.
@@
In writing the transition amplitude in the form of Eq. (66), one implicitly uses a participant/spectator approximation, assuming that the reaction occurs due to the interaction of the incident proton with a single neutron $(N_{1})$ of $A$ (the participant), leaving the residual $B=N_{2}+C$ system unperturbed.
To reduce Eq. (66) to a tractable form, one can approximate the exact wave function $\Psi^{(-)}_{f}$ by the factorized form,
@@
After integration over the angle of $\vec{q}$ and transforming from momentum to energy, expression (70) provides a double differential cross section with respect to the scattering angle and the excitation energy of the residual $N_{2}+C$ system. This observable can in principle be reconstructed experimentally from the measured energy and angle of the outgoing deuteron. To illustrate this formalism, we show in Fig. 32 results for the reaction 11Li(p,d)10Li Cas17 , recently measured at TRIUMF by Sanetullaev et al. San16 .
@@
For the calculations presented, optical potentials adjusted to describe $p,d+^{11}$i elastic scattering were used, following San16 . Besides, it was assumed that the two halo neutrons of 11Li have the same angular momentum $\vec{J}=\vec{L}+\vec{s}_{2}$. This condition is strictly fulfilled when the spin of the core is neglected and is also a good approximation when using $I=3/2^{-}$. Only the configurations $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ were considered, as other contributions, while present in the three-body calculations, give negligible components in the ground state of 11Li. The considered 9Li+n potential contains terms that couple the spins of the remaining neutron and 9Li in 10Li leading to splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ states and of the virtual state in the $s$-wave in $1^{-}$ and $2^{-}$ components. These components are shown in the top panel of Fig. 32, which depicts the relative energy spectrum between $n$ and 9Li. In red, the $p-$wave components are presented, where the splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ resonances at 0.37 and 0.61 MeV is evident, while in green the $s$-wave conributions are shown, presenting a virtual state (with scattering length $a=-37.9$ fm) only in the $2^{-}$ component. The total is shown in dashed blue, while its folding with the experimental resolution given in Ref. San16 is presented in the solid blue line. It is rather remarkable how the low-energy peaks are smoothed out by the experimental resolution leading to a relatively featureless distribution, which could explain why the fine details of 10Li are not observed in this experiment. In the bottom panel, the energy-integrated angular distribution of the outgoing deuteron is presented. It has been assumed that the subtraction of the non-resonant background performed in the analysis of the data completely excludes the $s$-wave component, so experimental data are compared only with the cross section corresponding to neutrons transferred from the $p_{1/2}$ orbitals, finding an excellent agreement both in shape and magnitude. Note that the experimental data have been measured at angles for which the $s_{1/2}$ contribution is minimal. This may explain why in the calculated energy distribution the contribution from $s$ waves is quite important at low energies, while experimentally there is no direct sign of this effect.
Using the methods described in Sec. 4.1 to consider continuum states of the $p-N$ subsystem, a similar formalism can be used to study $(p,pN)$ reactions on three-body nuclei, simply considering the auxiliary amplitude in Eq. (69) for continuum $p-N$ states, which are typically chosen to have a definite $p-N$ energy, angular momentum and parity:
@@
and where the outgoing $\chi_{pNB}^{(-)}$ functions can be computed using a CDCC solution for the final 3-body system Mor15 . Figure 33 shows results for the $n$-${}^{9}\mathrm{Li}$ relative energy distribution in the 11Li$(p,pn)^{10}$i reaction at 280 MeV/nucleon, using the same model for 11Li and 10Li as for the $(p,d)$ reaction in Fig. 32. The contributions of the different components are shown on the top panel, where again a splitting of the $p-$wave resonance is visible, as well as the virtual state only in the $2^{-}$ component, the main difference to the $(d,p)$ case being the increased importance of the $s$-wave components. In the bottom panel, the theoretical distribution is presented convoluted with the experimental resolution and compared with the experimental data, showing an excellent agreement with it, although the experimental resolution again significantly smooths the distribution.
It is a success of the model model that it is able to reproduce experimental data at $\sim$6 MeV per nucleon (the $(p,d)$ reaction) and at $\sim$280 MeV per nucleon (the $(p,pN)$ reaction), although perhaps not an unexpected one, given the similarities of both reactions when viewed as transfer to bound and continuum states of the $p-n$ system.